# Demonstration that Differential Length Changes of Optical Cavities are a Sensitive Probe for Ultralight Dark Matter

### v2

@@ Appendix A Isolated Sapphire Cavities at 6 K
Figure 6: Detailed optical layout of the cavities’ beat note and mechanical resonance measurement apparatus
@@
The fifth stage is a cryogenic 4-wire pendulum with a $\approx 1.5\,{\rm Hz}$ mechanical pole. The pendulum’s Molybdenum wires are $250\,\mu{\rm m}$ in diameter and 15 cm long. Each of the pendula supports an aluminum plate that, in turn, supports one optical cavity. The optical cavity spacers are supported by the aluminum platforms at 4 points that finite element calculations indicated would minimize vibration transmission [45, 46].
Heat is extracted from the cavities using high conductivity aluminum links that are 0.1 mm by 1 cm by 10 cm long. They are very pure and annealed to maximize the heat transfer, and to make them as flexible as possible to minimize vibration transmission. They are bonded to the sapphire using silver paint and clamped to the cavity platforms. Four-wire silicon diode sensors attached to the cavities with silver paint typically measure 6 K. At this temperature, we estimate a cavity finesse ($\mathcal{F}_{{\rm sapph}}$) of $\approx 120,000$.
Figure 7: The blue and orange dashed curves represent $A_{\rm det}$ (Eq. (7)) for the lowest and highest estimates, respectively, for the longitudinal resonances of the two cavities with the empirically estimated $Q_{{\scriptscriptstyle{\rm M}}}$. The black curve accounts for this uncertainty by conservatively taking the minimum of the two dashed curves.
@@ Appendix B Optics and Readout of Cavity Length
The laser beams after the sapphire cavities are recombined on a 50:50 non-polarizing beam splitter (NPBS) and detected with a 5 GHz beat note photodetector (PD, Thorlabs DET08C). The 262 MHz carrier is canceled by mixing with the same output after a delay line. A Fourier transform of the resulting time series from this “self-homodyne detection” is a spectrum as a function of the offset frequency from the carrier, insensitive to slow drifts of the carrier frequency. The discriminator is calibrated using the signal amplitude when an optical sideband is added with one of the AOMs.
Appendix C Detector response
@@
for a cavity pole frequency $f_{{\rm p}}$ [47, 46]. In our science band, $f_{\phi}\gg f_{{\rm p}}$ so that $H_{{\scriptscriptstyle{\rm OM}}}\approx 1$, making Eq. (4) a good approximation that is useful for an intuitive understanding.
There is uncertainty in our determination of the optical pole frequencies, and also of the mechanical resonance frequencies and quality factors of the two science cavities. The optical poles are estimated using $f_{{\rm p,}\,{\scriptscriptstyle{\rm L/S}}}=c/(4L_{{\scriptscriptstyle{\rm L/S}}}\mathcal{F}_{{\rm sapph}})$. The mechanical resonances at $f_{{\scriptscriptstyle{\rm M,\,L}}}\approx 34\,{\rm kHz}$ and $f_{{\scriptscriptstyle{\rm M,\,S}}}\approx 64\,{\rm kHz}$, are determined using a combination of finite element analysis and empirical verification by observing enhancement of peaks in the ASD resulting from a PZT drive. The mechanical quality factors are estimated by measuring the full-width at half maximum of the resonance peaks.
The uncertainty in our determination of the optical and mechanical pole frequencies of the two science cavities produces uncertainty in the detector response function. To be conservative, for the full detector response function we use the solid black curve in Fig. 7 in our analysis. This is the minimum of the two most extreme detector response functions (blue and orange dashed curves) we can make using the pole frequencies, resonance frequencies and quality factor values just described. In the future, improved characterization and optimization of the mechanical resonances would permit an improved search near the cavity resonance frequencies.
Appendix D Data Acquisition and Processing
@@
[…]
@@
×$\times\sinh\left(\eta\sqrt{\eta^{2}+4\pi(f_{p}-f_{\phi}^{\prime})\tau_{{\rm c}}}\right)\,,$ (11)
@@
($\mathcal{L}(\{\widetilde{d}_{p}\}|d_{m_{{\rm e}}})\equiv\prod_{p=1}^{N/2-1}\frac{1}{\pi\widetilde{\Sigma}_{p}}\exp\left[-\frac{|\widetilde{d}_{p}|^{2}}{\widetilde{\Sigma}_{p}}\right]\,,$ (12)
@@
We assume an uninformed prior, $p(d_{m_{{\rm e}}})$, and $p(\{\widetilde{d}_{p}\})$ is a normalization constant. For $N\Delta t\gg\tau_{\rm c}$, in an approximate analytic solution to Eq. (6), we expect the scaling $d_{m_{{\rm e}}}\propto\widetilde{h}/(N\Delta t\,\tau_{{\rm c}})^{1/4}$.

### v3

@@ Appendix A Isolated Sapphire Cavities at 6 K
The fifth stage is a cryogenic 4-wire pendulum with a $\approx 1.5\,{\rm Hz}$ mechanical pole. The pendulum’s Molybdenum wires are $250\,\mu{\rm m}$ in diameter and 15 cm long. Each pendulum supports an aluminum plate that, in turn, supports one optical cavity. The optical cavity spacers are supported by the aluminum platforms at 4 points that finite element calculations indicated would minimize vibration transmission [38, 39].
Heat is extracted from the cavities using high conductivity aluminum links (0.1 mm by 1 cm by 10 cm). They are very pure and annealed to maximize the heat transfer, and to make them as flexible as possible to minimize vibration transmission. They are bonded to the sapphire using silver paint and clamped to the cavity platforms. Four-wire silicon diode sensors attached to the cavities with silver paint typically measure 6 K.
Figure 7: Theoretical contributions of the thermal and shot noise for the sensitivity projections shown in Fig. 5.
@@ Appendix B Optics and Readout of Cavity Length
The laser beams after the sapphire cavities are recombined on a 50:50 non-polarizing beam splitter (NPBS) and detected with a 5 GHz beat note photodetector (PD) from Thorlabs (DET08C). The 262 MHz carrier is canceled by mixing with the same output after a delay line [40] and the resulting time series (i.e. $\delta f_{{\rm beat}}(t)$ from Eq. (3)) is stored on a data acquisition (DAQ) card. A Fourier transform of the time series from this “self-homodyne detection” corresponds to the optical strain spectrum (Fig. 2) as a function of the offset frequency from the carrier, insensitive to slow drifts of the carrier frequency. The discriminator is calibrated using the signal amplitude when an optical sideband is added with one of the AOMs (see [40] for calibration procedure). A summary of important apparatus parameters is shown in Table 1.
Description (notation) Values
Cavity lengths ($L_{{\scriptscriptstyle{\rm L/S}}}$) 15 cm / 7.5 cm
Mechanical resonances ($f_{{\scriptscriptstyle{\rm M,\,L/S}}}$) $34\,{\rm kHz}$ / $64\,{\rm kHz}$
Mechanical quality factor ($Q_{{\scriptscriptstyle{\rm M}}}$) $10^{4}$
Optical poles ($f_{{\rm p,}\,{\scriptscriptstyle{\rm L/S}}}$) $4.7\,{\rm kHz}$ / $7.4\,{\rm kHz}$
Optical frequency ($f_{\rm opt}$) $193\,{\rm THz}$
Input power per cavity 0.6 mW
Cavity power transmission efficiency 20%
Cavity finesse 120,000 ($\pm 12\%$)
Table 1: Summary of important apparatus parameter with nominal values.
Appendix C Theoretical limit estimation
The sensitivity projections in Fig. 5 were computed using the theoretical estimates of the cavities’ thermal and shot noise (Fig. 7). We assume our vibration isolation upgrades will be comparable to GW detectors where such noise is subdominant above $100\,{\rm Hz}$ [41]. Therefore, we conservatively ignore vibrational noise for $f_{\phi}>1\,{\rm kHz}$. Thermal noise is modeled by the formulas in [39]. The shot noise strain, $\widetilde{h}_{\rm shot}(f_{\phi})=f_{\phi}(2\pi\hbar(P_{{\scriptscriptstyle{\rm L}}}^{-1}+P_{{\scriptscriptstyle{\rm S}}}^{-1})/(\eta f_{\rm opt}))^{1/2}$, is determined by the powers ($P_{{\scriptscriptstyle{\rm L}}}$ and $P_{{\scriptscriptstyle{\rm S}}}$) from the long and short cavities incident on the photodetector, with quantum efficiency $\eta\approx 0.7$. $P_{{\scriptscriptstyle{\rm L}}}$ and $P_{{\scriptscriptstyle{\rm S}}}$ are balanced using a HWP before the PBS (Fig. 6). A second photodetector at the beam splitter’s other port can further improve the shot noise limit. Future increases in input power to 50 mW per cavity ($P_{{\scriptscriptstyle{\rm L/S}}}\approx 5\,{\rm mW}$ with 20% transmission) will not significantly increase thermal noise. Conservatively, assuming a cooling power no lower than now ($\approx(8\,{\rm mK}/{\rm mW})^{-1}$), we estimate a $\approx 0.4\,{\rm K}$ temperature increase.
Appendix D Detector response
@@
for a cavity pole frequency $f_{{\rm p}}$ [42, 39]. For $f_{\phi}\gg f_{{\rm p}}$, since $H_{{\scriptscriptstyle{\rm OM}}}\approx 1$, Eq. (4) is a good approximation that is useful for an intuitive understanding.
There is uncertainty in our determination of the optical pole frequencies, and also of the mechanical resonance frequencies and quality factors of the two science cavities. The optical poles ($f_{{\rm p,}\,{\scriptscriptstyle{\rm L/S}}}$) are estimated by fitting to the cavities’ measured optical transfer function. The mechanical resonances ($f_{{\scriptscriptstyle{\rm M,\,L/S}}}$) are empirically estimated by observing the enhancement of peaks in the ASD resulting from a PZT drive (Fig. 6). Mechanical quality factors ($Q_{{\scriptscriptstyle{\rm M}}}$) are estimated by measuring the full width at half maximum of the resonance peaks (Fig. 3 (b)). Measurements of $f_{{\scriptscriptstyle{\rm M}}}$ and $Q_{{\scriptscriptstyle{\rm M}}}$ serve as calibration for finite element analysis simulating a volumetric ULDM drive which agrees with the form of Eq. (5).
Appendix E Data Acquisition and Processing
@@
where $\widetilde{\phi}_{p}$ is the discrete Fourier transform of the ULDM field and $\langle\ldots\rangle$ denote a statistical average. Note that all of the above quantities are implicit functions of the ULDM Compton frequency $f_{\phi}$. Moreover, $\langle|\widetilde{\phi}_{p}|^{2}\rangle$ is defined in terms of the ULDM lineshape function $F_{p}$ as [23]
@@
$\times\sinh\left(\eta\sqrt{\eta^{2}+4\pi(f_{p}-f_{\phi}^{\prime})\tau_{{\rm c}}}\right)\,$, (11)
@@
$\mathcal{L}(\{\widetilde{d}_{p}\}|d_{m_{{\rm e}}})\equiv\prod_{p=1}^{N/2-1}\frac{1}{\pi\widetilde{\Sigma}_{p}}\exp\left[-\frac{|\widetilde{d}_{p}|^{2}}{\widetilde{\Sigma}_{p}}\right]\,$, (12)
@@
We assume an uninformed prior, $p(d_{m_{{\rm e}}})$, and $p(\{\widetilde{d}_{p}\})$ is a normalization constant.